OpenStudy (anonymous):

y varies directly with x. When y=12 x= 16. Find X when y =24

OpenStudy (anonymous):

x=32

OpenStudy (anonymous):

y~kx 12 = k*16 ---> k=12/16=3/4 so y = (3/4)x can you now find x when y=24?